Find the standard form of the equation of the specified hyperbola.
49x^2 - 196x - 36y^2 + 504y - 3332 = 0

Answer by lwsshak3(11628)

Find the standard form of the equation of the specified hyperbola.
49x^2 - 196x - 36y^2 + 504y - 3332 = 0
complete the square
49(x^2-4x+4)-36(y^2-14y+49)= 3332+196-1764
49(x-2)^2-36(y-7)^2=1764
divide by 1764
Equation of given hyperbola: (x-2)^2/36-(y-7)^2/49=1